\begin{thebibliography}{42}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Babenko and Lempitsky(2014)}]{babenko2014additive}
Artem Babenko and Victor Lempitsky. 2014.
\newblock Additive quantization for extreme vector compression.
\newblock In \emph{Proceedings of the IEEE Conference on Computer Vision and
  Pattern Recognition}, pages 931--938.

\bibitem[{Castellucci et~al.(2019)Castellucci, Bellomaria, Favalli, and
  Romagnoli}]{castellucci2019multi}
Giuseppe Castellucci, Valentina Bellomaria, Andrea Favalli, and Raniero
  Romagnoli. 2019.
\newblock Multi-lingual intent detection and slot filling in a joint bert-based
  model.
\newblock \emph{arXiv preprint arXiv:1907.02884}.

\bibitem[{Chen et~al.(2019)Chen, Zhuo, and Wang}]{chen2019bert}
Qian Chen, Zhu Zhuo, and Wen Wang. 2019.
\newblock Bert for joint intent classification and slot filling.
\newblock \emph{arXiv preprint arXiv:1902.10909}.

\bibitem[{Chen et~al.(2018)Chen, Min, and Sun}]{chen2018learning}
Ting Chen, Martin~Renqiang Min, and Yizhou Sun. 2018.
\newblock Learning k-way d-dimensional discrete codes for compact embedding
  representations.
\newblock \emph{arXiv preprint arXiv:1806.09464}.

\bibitem[{Chen and Sun(2019)}]{chen2019differentiable}
Ting Chen and Yizhou Sun. 2019.
\newblock Differentiable product quantization for end-to-end embedding
  compression.
\newblock \emph{arXiv preprint arXiv:1908.09756}.

\bibitem[{Chen et~al.(2016)Chen, Mou, Xu, Li, and Jin}]{chen2016}
Yunchuan Chen, Lili Mou, Yan Xu, Ge~Li, and Zhi Jin. 2016.
\newblock Compressing neural language models by sparse word representations.
\newblock \emph{arXiv preprint arXiv:1610.03950}.

\bibitem[{Coucke et~al.(2018)Coucke, Saade, Ball, Bluche, Caulier, Leroy,
  Doumouro, Gisselbrecht, Caltagirone, Lavril et~al.}]{coucke2018snips}
Alice Coucke, Alaa Saade, Adrien Ball, Th{\'e}odore Bluche, Alexandre Caulier,
  David Leroy, Cl{\'e}ment Doumouro, Thibault Gisselbrecht, Francesco
  Caltagirone, Thibaut Lavril, et~al. 2018.
\newblock Snips voice platform: an embedded spoken language understanding
  system for private-by-design voice interfaces.
\newblock \emph{arXiv preprint arXiv:1805.10190}.

\bibitem[{Devlin et~al.(2018)Devlin, Chang, Lee, and Toutanova}]{bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2018.
\newblock Bert: Pre-training of deep bidirectional transformers for language
  understanding.
\newblock \emph{arXiv preprint arXiv:1810.04805}.

\bibitem[{Einolghozati et~al.(2019)Einolghozati, Pasupat, Gupta, Shah, Mohit,
  Lewis, and Zettlemoyer}]{einolghozati2019improving}
Arash Einolghozati, Panupong Pasupat, Sonal Gupta, Rushin Shah, Mrinal Mohit,
  Mike Lewis, and Luke Zettlemoyer. 2019.
\newblock Improving semantic parsing for task oriented dialog.
\newblock \emph{arXiv preprint arXiv:1902.06000}.

\bibitem[{Faruqui et~al.(2015)Faruqui, Tsvetkov, Yogatama, Dyer, and
  Smith}]{faruqui}
Manaal Faruqui, Yulia Tsvetkov, Dani Yogatama, Chris Dyer, and Noah Smith.
  2015.
\newblock Sparse overcomplete word vector representations.
\newblock \emph{arXiv preprint arXiv:1506.02004}.

\bibitem[{Ganesh et~al.(2020)Ganesh, Chen, Lou, Khan, Yang, Chen, Winslett,
  Sajjad, and Nakov}]{ganesh2020compressing}
Prakhar Ganesh, Yao Chen, Xin Lou, Mohammad~Ali Khan, Yin Yang, Deming Chen,
  Marianne Winslett, Hassan Sajjad, and Preslav Nakov. 2020.
\newblock Compressing large-scale transformer-based models: A case study on
  bert.
\newblock \emph{arXiv preprint arXiv:2002.11985}.

\bibitem[{Goo et~al.(2018)Goo, Gao, Hsu, Huo, Chen, Hsu, and
  Chen}]{goo2018slot}
Chih-Wen Goo, Guang Gao, Yun-Kai Hsu, Chih-Li Huo, Tsung-Chieh Chen, Keng-Wei
  Hsu, and Yun-Nung Chen. 2018.
\newblock Slot-gated modeling for joint slot filling and intent prediction.
\newblock In \emph{Proceedings of the 2018 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 2 (Short Papers)}, pages 753--757.

\bibitem[{Guo et~al.(2014)Guo, Tur, Yih, and Zweig}]{guo2014joint}
Daniel Guo, Gokhan Tur, Wen-tau Yih, and Geoffrey Zweig. 2014.
\newblock Joint semantic utterance classification and slot filling with
  recursive neural networks.
\newblock In \emph{2014 IEEE Spoken Language Technology Workshop (SLT)}, pages
  554--559. IEEE.

\bibitem[{Gupta et~al.(2018)Gupta, Shah, Mohit, Kumar, and
  Lewis}]{gupta2018semantic}
Sonal Gupta, Rushin Shah, Mrinal Mohit, Anuj Kumar, and Mike Lewis. 2018.
\newblock Semantic parsing for task oriented dialog using hierarchical
  representations.
\newblock \emph{arXiv preprint arXiv:1810.07942}.

\bibitem[{Hakkani-T{\"u}r et~al.(2016)Hakkani-T{\"u}r, T{\"u}r, Celikyilmaz,
  Chen, Gao, Deng, and Wang}]{hakkani2016multi}
Dilek Hakkani-T{\"u}r, G{\"o}khan T{\"u}r, Asli Celikyilmaz, Yun-Nung Chen,
  Jianfeng Gao, Li~Deng, and Ye-Yi Wang. 2016.
\newblock Multi-domain joint semantic frame parsing using bi-directional
  rnn-lstm.
\newblock In \emph{Interspeech}, pages 715--719.

\bibitem[{Jang et~al.(2016)Jang, Gu, and Poole}]{jang2016categorical}
Eric Jang, Shixiang Gu, and Ben Poole. 2016.
\newblock Categorical reparameterization with gumbel-softmax.
\newblock \emph{arXiv preprint arXiv:1611.01144}.

\bibitem[{Kamath and Das(2018)}]{kamath2018survey}
Aishwarya Kamath and Rajarshi Das. 2018.
\newblock A survey on semantic parsing.
\newblock \emph{arXiv preprint arXiv:1812.00978}.

\bibitem[{Kim(2014)}]{kim2014}
Yoon Kim. 2014.
\newblock Convolutional neural networks for sentence classification.
\newblock \emph{CoRR}, abs/1408.5882.

\bibitem[{Lai et~al.(2017)Lai, Xie, Liu, Yang, and Hovy}]{lai2017race}
Guokun Lai, Qizhe Xie, Hanxiao Liu, Yiming Yang, and Eduard Hovy. 2017.
\newblock Race: Large-scale reading comprehension dataset from examinations.
\newblock \emph{arXiv preprint arXiv:1704.04683}.

\bibitem[{Lam(2018)}]{word2bits}
Maximilian Lam. 2018.
\newblock Word2bits-quantized word vectors.
\newblock \emph{arXiv preprint arXiv:1803.05651}.

\bibitem[{Lan et~al.(2019)Lan, Chen, Goodman, Gimpel, Sharma, and
  Soricut}]{albert}
Zhenzhong Lan, Mingda Chen, Sebastian Goodman, Kevin Gimpel, Piyush Sharma, and
  Radu Soricut. 2019.
\newblock Albert: A lite bert for self-supervised learning of language
  representations.
\newblock \emph{arXiv preprint arXiv:1909.11942}.

\bibitem[{Liu and Lane(2016)}]{liu2016attention}
Bing Liu and Ian Lane. 2016.
\newblock Attention-based recurrent neural network models for joint intent
  detection and slot filling.
\newblock \emph{arXiv preprint arXiv:1609.01454}.

\bibitem[{Liu et~al.(2019)Liu, Ott, Goyal, Du, Joshi, Chen, Levy, Lewis,
  Zettlemoyer, and Stoyanov}]{liu2019roberta}
Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer
  Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019.
\newblock Roberta: A robustly optimized bert pretraining approach.
\newblock \emph{arXiv preprint arXiv:1907.11692}.

\bibitem[{Loshchilov and Hutter(2017)}]{loshchilov2017decoupled}
Ilya Loshchilov and Frank Hutter. 2017.
\newblock Decoupled weight decay regularization.
\newblock \emph{arXiv preprint arXiv:1711.05101}.

\bibitem[{Maddison et~al.(2016)Maddison, Mnih, and Teh}]{maddison2016concrete}
Chris~J Maddison, Andriy Mnih, and Yee~Whye Teh. 2016.
\newblock The concrete distribution: A continuous relaxation of discrete random
  variables.
\newblock \emph{arXiv preprint arXiv:1611.00712}.

\bibitem[{Mesnil et~al.(2013)Mesnil, He, Deng, and
  Bengio}]{mesnil2013investigation}
Gr{\'e}goire Mesnil, Xiaodong He, Li~Deng, and Yoshua Bengio. 2013.
\newblock Investigation of recurrent-neural-network architectures and learning
  methods for spoken language understanding.
\newblock In \emph{Interspeech}, pages 3771--3775.

\bibitem[{Mikolov et~al.(2013)Mikolov, Chen, Corrado, and
  Dean}]{mikolov2013efficient}
Tomas Mikolov, Kai Chen, Greg Corrado, and Jeffrey Dean. 2013.
\newblock Efficient estimation of word representations in vector space.
\newblock \emph{arXiv preprint arXiv:1301.3781}.

\bibitem[{Pennington et~al.(2014)Pennington, Socher, and
  Manning}]{pennington2014glove}
Jeffrey Pennington, Richard Socher, and Christopher~D Manning. 2014.
\newblock Glove: Global vectors for word representation.
\newblock In \emph{Proceedings of the 2014 conference on empirical methods in
  natural language processing (EMNLP)}, pages 1532--1543.

\bibitem[{Price(1990)}]{price1990evaluation}
Patti Price. 1990.
\newblock Evaluation of spoken language systems: The atis domain.
\newblock In \emph{Speech and Natural Language: Proceedings of a Workshop Held
  at Hidden Valley, Pennsylvania, June 24-27, 1990}.

\bibitem[{Rajpurkar et~al.(2016)Rajpurkar, Zhang, Lopyrev, and
  Liang}]{rajpurkar2016squad}
Pranav Rajpurkar, Jian Zhang, Konstantin Lopyrev, and Percy Liang. 2016.
\newblock Squad: 100,000+ questions for machine comprehension of text.
\newblock \emph{arXiv preprint arXiv:1606.05250}.

\bibitem[{Raunak(2017)}]{raunak2017}
Vikas Raunak. 2017.
\newblock Simple and effective dimensionality reduction for word embeddings.
\newblock \emph{arXiv preprint arXiv:1708.03629}.

\bibitem[{Rongali et~al.(2020)Rongali, Soldaini, Monti, and Hamza}]{rongali}
Subendhu Rongali, Luca Soldaini, Emilio Monti, and Wael Hamza. 2020.
\newblock Don't parse, generate! a sequence to sequence architecture for
  task-oriented semantic parsing.
\newblock \emph{arXiv preprint arXiv:2001.11458}.

\bibitem[{Sabour et~al.(2017)Sabour, Frosst, and Hinton}]{sabour2017}
Sara Sabour, Nicholas Frosst, and Geoffrey~E. Hinton. 2017.
\newblock Dynamic routing between capsules.
\newblock \emph{CoRR}, abs/1710.09829.

\bibitem[{Sanh et~al.(2019)Sanh, Debut, Chaumond, and Wolf}]{distilbert}
Victor Sanh, Lysandre Debut, Julien Chaumond, and Thomas Wolf. 2019.
\newblock Distilbert, a distilled version of bert: smaller, faster, cheaper and
  lighter.
\newblock \emph{arXiv preprint arXiv:1910.01108}.

\bibitem[{Shu and Nakayama(2017)}]{codebooks}
Raphael Shu and Hideki Nakayama. 2017.
\newblock Compressing word embeddings via deep compositional code learning.
\newblock \emph{arXiv preprint arXiv:1711.01068}.

\bibitem[{Subramanian et~al.(2018)Subramanian, Pruthi, Jhamtani,
  Berg-Kirkpatrick, and Hovy}]{spine}
Anant Subramanian, Danish Pruthi, Harsh Jhamtani, Taylor Berg-Kirkpatrick, and
  Eduard Hovy. 2018.
\newblock Spine: Sparse interpretable neural embeddings.
\newblock In \emph{Thirty-Second AAAI Conference on Artificial Intelligence}.

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones,
  Gomez, Kaiser, and Polosukhin}]{attention}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, {\L}ukasz Kaiser, and Illia Polosukhin. 2017.
\newblock Attention is all you need.
\newblock In \emph{Advances in neural information processing systems}, pages
  5998--6008.

\bibitem[{Vinyals et~al.(2015)Vinyals, Fortunato, and
  Jaitly}]{vinyals2015pointer}
Oriol Vinyals, Meire Fortunato, and Navdeep Jaitly. 2015.
\newblock Pointer networks.
\newblock In \emph{Advances in neural information processing systems}, pages
  2692--2700.

\bibitem[{Wang et~al.(2018)Wang, Singh, Michael, Hill, Levy, and
  Bowman}]{GLUEdataset}
Alex Wang, Amanpreet Singh, Julian Michael, Felix Hill, Omer Levy, and
  Samuel~R. Bowman. 2018.
\newblock {GLUE:} {A} multi-task benchmark and analysis platform for natural
  language understanding.
\newblock \emph{CoRR}, abs/1804.07461.

\bibitem[{Wolf et~al.(2019)Wolf, Debut, Sanh, Chaumond, Delangue, Moi, Cistac,
  Rault, Louf, Funtowicz, and Brew}]{Wolf2019HuggingFacesTS}
Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue,
  Anthony Moi, Pierric Cistac, Tim Rault, R'emi Louf, Morgan Funtowicz, and
  Jamie Brew. 2019.
\newblock Huggingface's transformers: State-of-the-art natural language
  processing.
\newblock \emph{ArXiv}, abs/1910.03771.

\bibitem[{Xu and Sarikaya(2013)}]{xu2013convolutional}
Puyang Xu and Ruhi Sarikaya. 2013.
\newblock Convolutional neural network based triangular crf for joint intent
  detection and slot filling.
\newblock In \emph{2013 ieee workshop on automatic speech recognition and
  understanding}, pages 78--83. IEEE.

\bibitem[{Zhang et~al.(2019)Zhang, Li, Du, Fan, and Yu}]{zhang2019}
Chenwei Zhang, Yaliang Li, Nan Du, Wei Fan, and Philip Yu. 2019.
\newblock Joint slot filling and intent detection via capsule neural networks.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, Florence, Italy. Association for
  Computational Linguistics.

\end{thebibliography}

\pdfoutput=1
\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{emnlp2020}
\usepackage{times}
\usepackage{latexsym}
\usepackage{graphicx, subfigure, multirow}
\usepackage{amsmath}
\usepackage{bbm}
\usepackage{amsfonts}

\newcommand{\argmin}{\mathop{\arg\min}}

\usepackage{microtype}

\aclfinalcopy 
\def\aclpaperid{3476} 

\title{Compressing Transformer-Based Semantic Parsing Models using Compositional Code Embeddings}

\author{Prafull Prakash$^{1} \thanks{~~Equal contribution, alphabetical order}$~, Saurabh Kumar Shashidhar$^{1}$\footnotemark[1]~, Wenlong Zhao$^{1}$\footnotemark[1]~, \\
{\bf Subendhu Rongali$^1$, Haidar Khan$^2$, \and Michael Kayser$^2$} \\
  $^1$University of Massachusetts Amherst \\
  $^2$Amazon Alexa \\
  \texttt{\{prafullpraka, ssaurabhkuma, wenlongzhao, srongali\}@cs.umass.edu}\\
  \texttt{\{khhaida, mikayser\}@amazon.com}}

\begin{document}
\maketitle
\begin{abstract}
The current state-of-the-art task-oriented semantic parsing models use BERT or RoBERTa as pretrained encoders; these models have huge memory footprints. This poses a challenge to their deployment for voice assistants such as Amazon Alexa and Google Assistant on edge devices with limited memory budgets. We propose to learn compositional code embeddings to greatly reduce the sizes of BERT-base and RoBERTa-base. We also apply the technique to DistilBERT, ALBERT-base, and ALBERT-large, three already compressed BERT variants which attain similar state-of-the-art performances on semantic parsing with much smaller model sizes. We observe 95.15\% $\sim$ 98.46\% embedding compression rates and 20.47\% $\sim$ 34.22\% encoder compression rates, while preserving $>$97.5\% semantic parsing performances. We provide the recipe for training and analyze the trade-off between code embedding sizes and downstream performances.
\end{abstract}

\section{Introduction}
Conversational virtual assistants, such as Amazon Alexa, Google Home, and Apple Siri, have become increasingly popular in recent times. These systems can process queries from users and perform tasks such as playing music and finding locations. A core component in these systems is a task-oriented semantic parsing model that maps natural language expressions to structured representations containing intents and slots that describe the task to perform. For example, the expression \emph{Can you play some songs by Coldplay?} may be converted to \emph{Intent: PlaySong, Artist: Coldplay}, and the expression \emph{Turn off the bedroom light} may be converted to \emph{Intent: TurnOffLight, Device: bedroom}. 

Task-oriented semantic parsing is traditionally approached as a joint intent classification and slot filling task. \citet{kamath2018survey} provide a comprehensive survey of models proposed to solve this task. Researchers have developed semantic parsers based on Recurrent Neural Networks \cite{mesnil2013investigation, liu2016attention, hakkani2016multi}, Convolutional Neural Networks \cite{xu2013convolutional, kim2014}, Recursive Neural Networks \cite{guo2014joint}, Capsule Networks \cite{sabour2017, zhang2019}, and slot-gated attention-based models \cite{goo2018slot}. 

The current state-of-the-art models on SNIPS \cite{coucke2018snips}, ATIS \cite{price1990evaluation}, and Facebook TOP \cite{gupta2018semantic} datasets are all based on BERT-style \cite{bert, liu2019roberta} encoders and transformer architectures \cite{chen2019bert,castellucci2019multi,rongali}. It is challenging to deploy these large models on edge devices and enable the voice assistants to operate locally instead of relying on central cloud services, due to the limited memory budgets on these devices. However, there has been a growing push towards the idea of TinyAI \footnote{https://www.technologyreview.com/technology/tiny-ai/}.

In this paper, we aim to build space-efficient task-oriented semantic parsing models that produce near state-of-the-art performances by compressing existing large models. We propose to learn compositional code embeddings to significantly compress BERT-base and RoBERTa-base encoders with little performance loss. We further use ALBERT-base/large \cite{albert} and DistilBERT \cite{distilbert} to establish light baselines that achieve similar state-of-the-art performances, and apply the same code embedding technique. We show that our technique is complementary to the compression techniques used in ALBERT and DistilBERT. With all variants, we achieve 95.15\% $\sim$ 98.46\% embedding compression rates and 20.47\% $\sim$ 34.22\% encoder compression rates, with $>$97.5\% semantic parsing performance preservation.

\section{Related Compression Techniques}
\subsection{BERT Compression}
Many techniques have been proposed to compress BERT \cite{bert}. \citet{ganesh2020compressing} provide a survey on these methods. Most existing methods focus on alternative architectures in transformer layers or learning strategies. 

In our work, we use DistilBERT and ALBERT-base as light pretrained language model encoders for semantic parsing. DistilBERT \cite{distilbert} uses distillation to pretrain a model that is 40\% smaller and 60\% faster than BERT-base, while retaining 97\% of its downstream performances. ALBERT \cite{albert} factorizes the embedding and shares parameters among the transformer layers in BERT and results in better scalability than BERT. ALBERT-xxlarge outperforms BERT-large on GLUE \cite{GLUEdataset}, RACE \cite{lai2017race}, and SQUAD \cite{rajpurkar2016squad} while using less parameters.

We use compositional code learning \cite{codebooks} to compress the model embeddings, which contain a substantial amount of model parameters. Previously ALBERT uses factorization to compress the embeddings. We find more compression possible with code embeddings.

\subsection{Embedding Compression}
Varied techniques have been proposed to learn compressed versions of non-contextualized word embeddings, such as, Word2Vec \cite{mikolov2013efficient} and GLoVE \cite{pennington2014glove}. \citet{spine} use denoising k-sparse autoencoders to achieve binary sparse intrepretable word embeddings. \citet{chen2016} achieve sparsity by representing the embeddings of uncommon words using sparse linear common combination of common words. \citet{word2bits} achieve compression by quantization of the word embeddings by using 1-2 bits per parameter. \citet{faruqui} use sparse coding in a dictionary learning setting to obtain sparse, non-negative word embeddings. \citet{raunak2017} achieve dense compression of word embeddings using PCA combined with a post-processing algorithm. \citet{codebooks} propose to represent word embeddings using compositional codes learnt directly in end-to-end fashion using neural networks. Essentially few common basis vectors are learnt and embeddings are reconstructed using their composition via a discrete code vector specific to each token embedding. This results in 98\% compression rate in sentiment analysis and 94\% - 99\% in machine translation tasks without performance loss with LSTM based models. All the above techniques are applied to embeddings such as WordVec and Glove, or LSTM models. 

We aim to learn space-efficient embeddings for transformer-based models. We focus on compositional code embeddings \cite{codebooks} since they maintain the vector dimensions, do not require special kernels for calculating in a sparse or quantized space, can be finetuned with transformer-based models end-to-end, and achieve extremely high compression rate. \citet{chen2018learning} explores similar idea as \citet{codebooks} and experiment with more complex composition functions and guidances for training the discrete codes. \citet{chen2019differentiable} further show that end-to-end training from scratch of models with code embeddings is possible. Given various pretrained language models, we find that the method proposed by \citet{codebooks} is straightforward and perform well in our semantic parsing experiments.

\section{Method}
\subsection{Compositional Code Embeddings}
\citet{codebooks} apply additive quantization \cite{babenko2014additive} to learn compositional code embeddings to reconstruct pretrained word embeddings such as GloVe \cite{pennington2014glove}, or task-specific model embeddings such as those from an LSTM neural machine translation model. Compositional code embeddings $E^C$ for vocabulary $V$ consist of a set of $M$ codebooks $E_1^C$, $E_2^C$, ..., $E_M^C$, each with $K$ basis vectors of the same dimensionality $D$ as the reference embeddings $E$, and a discrete code vector ($C_w^1$, $C_w^2$, ..., $C_w^M$) for each token $w$ in the vocabulary. The final embedding for $w$ is composed by summing up the $C_w^i$th vector from the ith codebook as $E^C(C_w)=\sum_{i=1}^M E_i^C(C_w^i)$. Codebooks and discrete codes are jointly learned using the mean squared distance objective: $(C^*, E^{C*})=\argmin_{C, E^C} \frac{1}{|V|}\sum_{w\in V}||E^C(C_w)-E(w)||^2.$ For learning compositional codes, the Gumbel-softmax reparameterization trick \cite{jang2016categorical, maddison2016concrete} is used for one-hot vectors corresponding to each discrete code.

\begin{table*}[t]
\centering
\resizebox{\textwidth}{!}{%
\begin{tabular}{lcccccccc}
\hline \textbf{Encoder} & \textbf{EncoderParam\# / Size} &\textbf{EmbParam\# / Size} & \textbf{SizeRatio}& \textbf{CCEmbSize} & \textbf{CCEncoderSize} & \textbf{EmbComp} & \textbf{EncoderComp}\\ \hline
RoBERTa-base            & 125.29M / 477.94MB  & 38.60M / 147.25MB   & 30.81\%   & 2.27MB    & 332.96MB & 98.46\%   & 30.33\% \\
BERT-base-uncased	    & 110.10M / 420.00MB  & 23.44M / 89.42MB    & 21.29\%   & 1.97MB    & 332.55MB & 97.80\%   & 20.82\% \\
DistilBERT-base-uncased	& 66.99M / 255.55MB   & 23.44M / 89.42MB    & 34.99\%   & 1.97MB    & 168.10MB & 97.80\%   & 34.22\% \\
ALBERT-large-v2         & 17.85M / 68.09MB    & 3.84M / 14.65MB     & 21.52\%   & 0.71MB    & 54.15MB  & 95.15\%   & 20.47\% \\
ALBERT-base-v2          & 11.81M / 45.05MB    & 3.84M / 14.65MB     & 32.52\%   & 0.71MB    & 31.11MB  & 95.15\%   & 30.94\% \\
\hline
\end{tabular}
}
\caption{\label{adv} Model compression with compositional code (``cc") embeddings. The embedding layers are compressed by more than 95\% with compositional code embeddings in all of the BERT variants.}
\end{table*}

\subsection{Transformer-Based Models with Compositional Code Embeddings}
In this work, we learn compositional code embeddings to reduce the size of the embeddings in pretrained contextualized language models. We extract the embedding tables from pretrained RoBERTa-base \cite{liu2019roberta}, BERT-base \cite{bert}, DistilBERT-base \cite{distilbert}, ALBERT-large-v2 and ALBERT-base-v2 \cite{albert} from the huggingface transformers library \cite{Wolf2019HuggingFacesTS} and follow the approach presented by \citet{codebooks} to learn the code embeddings. We then replace the embedding tables in the transformer models with the compositional code approximations and evaluate the compressed language models by finetuning on downstream tasks. When \citet{codebooks} feed compositional code embeddings into the LSTM neural machine translation model, they fix the embedding parameters and train the rest of the model from random initial values. In our experiments, we fix the discrete codes, initialize the transformer layers with those from the pretrained language models, initialize the task-specific output layers randomly, and finetune the codebook basis vectors with the rest of the non-discrete parameters.

\subsection{Size Advantage of Compositional Code Embeddings}
An embedding matrix $E\in \mathbb{R}^{|V|\times D}$ stored as 32-bit float point numbers, where $|V|$ is the vocabulary size and $D$ is the embedding dimension, requires $32|V|D$ bits. Its compositional code reconstruction requires $32MKD$ bits for $MK$ basis vectors, and $M\log_2K$ bits for codes of each of $|V|$ tokens. Since each discrete code takes an integer value in $[1, K]$, it can be represented using $\log_2K$ bits. 

Table \ref{adv} illustrates the size advantage of compositional code embeddings for various pretrained transformer models \cite{Wolf2019HuggingFacesTS} used in our experiments. While the technique focuses on compressing the embedding table, it is compatible with other compression techniques for transformer models, including parameter sharing among transformer layers and embedding factorization used in ALBERT and distillation for learning DistilBERT. In our experiments, we apply the code learning technique to compress embeddings in five pretrained BERT variants by 95.15\% $\sim$ 98.46\% to build competitive but significantly lighter semantic parsing models.

\begin{table}
\centering
\resizebox{0.45\textwidth}{!}{%
\begin{tabular}{lccccc}
\hline \textbf{Dataset} & \textbf{Train} &\textbf{Valid} & \textbf{Test} & \textbf{\#Intent} & \textbf{\#Slot} \\ \hline
ATIS	& 4,478 & 500 & 893 & 26 & 83 \\
SNIPS	& 13,084 & 700 & 700 & 7 & 39 \\
Facebook TOP    & 31,279 & 4,462 & 9,042 & 25 & 36 \\
\hline
\end{tabular}
}
\caption{\label{table-data} Statistics for semantic parsing datasets.}
\end{table}

\section{Datasets}
Following \citet{rongali}, we evaluate our models on SNIPS \cite{coucke2018snips}, Airline Travel Information
System (ATIS) \cite{price1990evaluation}, and Facebook TOP \cite{gupta2018semantic} datasets for task-oriented semantic parsing (Table \ref{table-data}). For SNIPS and ATIS, we use the same train/validation/test split as \citet{goo2018slot}.

\section{Experiments and Analyses}
For transformer model training, we base our implementation on the huggingface transformers library v2.6.0 \cite{Wolf2019HuggingFacesTS}. We use the AdamW optimizer \cite{loshchilov2017decoupled} with 10\% warmup steps and linear learning rate decay to 0. Forr code embedding learning, we base our implementation on that of \citet{codebooks}. By default we learn code embeddings with 32 codebooks and 16 basis vectors per codebook. Unless otherwise specified, hyperparameters are found according to validation performances from one random run. We conduct our experiments on a mixture of Tesla M40, TITAN X, 1080 Ti, and 2080 Ti GPUs. We use exact match (EM) and intent accuracy as evaluation metrics. Exact match requires correct predictions for all intents and slots in a query, and is our primary metric.

\begin{table}[t]
\centering
\resizebox{0.5\textwidth}{!}{%
\begin{tabular}{lcccccc}
\hline \multicolumn{5}{l}{\textbf{Model}} & \textbf{EM} &\textbf{Intent}  \\ \hline
\multicolumn{5}{l}{Joint BiRNN \cite{hakkani2016multi}}	& 73.2 & 96.9 \\
\multicolumn{5}{l}{Attention BiRNN \cite{liu2016attention}}	&  74.1 & 96.7 \\
\multicolumn{5}{l}{Slot Gated Full Attention \cite{goo2018slot}}	& 75.5 & 97.0 \\
\multicolumn{5}{l}{CapsuleNLU \cite{zhang2019}}	&  80.9 & 97.3 \\
\multicolumn{5}{l}{BERT-Seq2Seq-Ptr \cite{rongali}}	& 86.3 & 98.3 \\
\multicolumn{5}{l}{RoBERTa-Seq2Seq-Ptr \cite{rongali}}	& 87.1 & 98.0 \\
\multicolumn{5}{l}{BERT-Joint \cite{castellucci2019multi}}	& 91.6 & \textbf{99.0} \\
\multicolumn{5}{l}{Joint BERT \cite{chen2019bert}}	& \textbf{92.8} & 98.6  \\
\hline \textbf{Ours} & \textbf{epo} & \textbf{lr} & \textbf{wd} & \textbf{EM-v} &\textbf{EM} & \textbf{Intent} \\ \hline
ALBERT-base	&   & 5e-5 & 0.05 & 90.71 & \textbf{91.29} & 98.86 \\
ALBERT-base\_cc & 1100 & 5e-5 & 0.01 & 90.00 & 89.14 & 98.14 \\\hline
ALBERT-large	&   & 3e-5 & 0.05 & 91.29 & \textbf{92.43} & 98.14 \\
ALBERT-large\_cc & 1100 & 2e-5 & 0.05 & 91.14 & \textbf{92.43} & 98.71 \\\hline
DistilBERT-base	&   & 3e-5 & 0.05 & 90.29 & 91.14 & 98.57 \\
DistilBERT-base\_cc	& 900 & 6e-5 & 0.01 & 90.14 & \textbf{91.24} & 98.43 \\\hline
BERT-base	&  & 3e-5 & 0.05 & 92.14 & \textbf{92.29} & 99.14  \\
BERT-base\_cc	& 900 & 6e-5  & 0.05 & 91.29 & 90.71 & 98.71 \\
\hline
\end{tabular}
}
\caption{\label{table-SNIPS} Results on SNIPS. ``cc" indicate models with code embeddings. ``epo" is the epoch number for offline code embedding learning. ``lr" and ``wd" are the peak learning rate and weight decay for whole model finetuning. ``EM-v", ``EM", ``Intent" indicate validation exact match, test exact match, and test intent accuracy.}
\end{table}

\subsection{SNIPS and ATIS}
We implement a joint sequence-level and token-level classification layer for pretrained transformer models. The intent probabilities are predicted as $\mathrm{y^i=softmax(W^ih_0+b^i)}$, where $\mathrm{h_0}$ is the hidden state of the [CLS] token. The slot probabilities for each token j are predicted as $\mathrm{y^s_j=softmax(W^sh_j+b^s)}$. We use the cross entropy loss to maximize $\mathrm{p(y^i|x)\prod p(y^s_j|x)}$ where j is the first piece-wise token for each word in the query. We learn code embeddings for \{500, 700, 900, 1100, 1300\} epochs. We train transformer models with original and code embeddings all for 40 epochs with batch size 16 and sequence length 128. Uncased BERT and DistilBERT perform better than the cased versions. We experiment with peak learning rate \{2e-5, 3e-5, ..., 6e-5\} and weight decay \{0.01, 0.05, 0.1\}. As shown in Table \ref{table-SNIPS} and \ref{table-ATIS}, we use different transformer encoders to establish strong baselines which achieve EM values that are within 1.5\% of the state-of-the-art.

On both datasets, models based on our compressed ALBERT-large-v2 encoder (54MB) perserves $>$99.6\% EM of the previous state-of-the-art model \cite{chen2019bert} which uses a BERT encoder (420MB). In all settings, our compressed encoders perserve $>$97.5\% EM of the uncompressed counterparts under the same training settings. We show that our technique is effective on a variety of pretrained transformer encoders.

\begin{table}[t]
\centering
\resizebox{0.5\textwidth}{!}{%
\begin{tabular}{lcccccc}
\hline \multicolumn{5}{l}{\textbf{Model}} & \textbf{EM} &\textbf{Intent}  \\ \hline
\multicolumn{5}{l}{Joint-BiRNN \cite{hakkani2016multi}}	& 80.7 & 92.6\\
\multicolumn{5}{l}{Attention-BiRNN \cite{liu2016attention}}	& 78.9 & 91.1\\
\multicolumn{5}{l}{Slot-Gated \cite{goo2018slot}}	& 82.2 & 93.6\\
\multicolumn{5}{l}{CapsuleNLU \cite{zhang2019}}	& 83.4 & 95.0\\
\multicolumn{5}{l}{BERT-Seq2Seq-Ptr \cite{rongali}}	& 86.4 & 97.4 \\
\multicolumn{5}{l}{RoBERTa-Seq2Seq-Ptr \cite{rongali}}	& 87.1 & 97.4\\
\multicolumn{5}{l}{BERT-Joint \cite{castellucci2019multi}}	& \textbf{88.2} & \textbf{97.8} \\
\multicolumn{5}{l}{Joint-BERT \cite{chen2019bert}}	& \textbf{88.2} & 97.5 \\ 
\hline \textbf{Ours} & \textbf{epo} & \textbf{lr} & \textbf{wd} & \textbf{EM-v} &\textbf{EM} & \textbf{Intent} \\ \hline
ALBERT-base	&   & 5e-5 & 0.05 & 93.4 & 86.90 & 97.42 \\
ALBERT-base\_cc & 900 & 6e-5 & 0.1 & 94.2 & \textbf{87.23} & 96.75 \\\hline
ALBERT-large	&   & 5e-5 & 0.05 & 93.8 & \textbf{88.02} & 97.54 \\
ALBERT-large\_cc & 1100 & 5e-5 & 0.1 & 94.0 & 87.91 & 97.54 \\\hline
DistilBERT-base	&   & 4e-5 & 0.05 & 93.6 & \textbf{88.13} & 97.42 \\
DistilBERT-base\_cc	& 1100 & 6e-5 & 0.05 & 93.2 & 87.12 & 97.54 \\\hline
BERT-base	&   &  4e-5 & 0.01 & 93.4 & \textbf{88.13} & 97.54 \\
BERT-base\_cc	& 700 & 6e-5  & 0.1 & 93.0 & 87.35 & 97.20 \\
\hline
\end{tabular}
}
\caption{\label{table-ATIS} Results on ATIS. Refer to the caption of Table \ref{table-SNIPS} for abbreviation explanations.}
\end{table}

\subsection{Facebook TOP}
\begin{table}[t]
\centering
\resizebox{0.5\textwidth}{!}{%
\begin{tabular}{lccc}
\hline \multicolumn{2}{l}{\textbf{Model}} & \textbf{EM} &\textbf{Intent}  \\ \hline
\multicolumn{2}{l}{RNNG \cite{gupta2018semantic}} & 78.51 & - \\ 
\multicolumn{2}{l}{Shift Reduce (SR) Parser} & 80.86 & - \\
\multicolumn{2}{l}{SR with ELMo embeddings} &83.93 & - \\
\multicolumn{2}{l}{SR ensemble + ELMo + SVMRank} & \textbf{87.25} & - \\
\multicolumn{2}{l}{BERT-Seq2Seq-Ptr \cite{rongali}}	& 83.13 & 97.91 \\
\multicolumn{2}{l}{RoBERTa-Seq2Seq-Ptr \cite{rongali}}	& 86.67 & \textbf{98.13} \\
\hline \textbf{Ours} & \textbf{EM-v} &\textbf{EM} & \textbf{Intent} \\ \hline
ALBERT-Seq2Seq-Ptr & 84.56 & \textbf{85.41} & 98.47 \\
ALBERT-Seq2Seq-Ptr\_cc & 83.48 & 84.42 & 98.05 \\ \hline
DistilBERT-Seq2Seq-Ptr & 84.25 & \textbf{85.12} & 98.50 \\
DistilBERT-Seq2Seq-Ptr\_cc & 82.76 & 83.42 & 98.09 \\ \hline
BERT-Seq2Seq-Ptr & 83.83 & \textbf{85.01} & 98.59 \\
BERT-Seq2Seq-Ptr\_cc & 82.36 & 83.34 & 98.25 \\ \hline
RoBERTa-Seq2Seq-Ptr	 & 85.00 & \textbf{85.67} & 98.59 \\
RoBERTa-Seq2Seq-Ptr\_cc	& 83.51 & 83.78 & 98.17 \\ \hline
\end{tabular}
}
\caption{\label{table-BaselineTOP} Results on Facebook TOP. The SR models are by \citet{einolghozati2019improving}. Refer to the caption of Table \ref{table-SNIPS} for abbreviation explanations.}
\end{table}

Table \ref{table-BaselineTOP} presents results on Facebook TOP. We follow \citet{rongali} and experiment with Seq2Seq models. We use different pretrained BERT-variants as the encoder, transformer decoder layers with $d_{model}=768$ \cite{attention}, and a pointer generator network \cite{vinyals2015pointer} which uses scaled dot-product attention to score tokens. The model is trained using the cross-entropy loss with label smoothing of 0.1. For simplicity, we always train code embeddings for 900 epochs offline. Learning rate 2e-5 and weight decay 0.01 are used for transformer training. BERT and DistilBERT are cased in these experiments. During inference, we employ beam decoding with width 5. Our greatly compressed models present 98$\sim$99\% performances of the original models.

\begin{table*}
\centering
\tiny
\resizebox{.75\textwidth}{!}{%
\begin{tabular}{ccccccc}
\hline \textbf{Epoch} & \textbf{MeanEucDist} &\textbf{NN-cos} & \textbf{NN-Euc} & \textbf{SNIPS} & \textbf{ATIS} & \textbf{TOP}\\ \hline
100	    & 0.3677$\pm$0.25\%  & 0.66$\pm$1.90\%        & 0.65$\pm$2.00\% & 79.29 & 82.31 & 78.09 \\
200	    & 0.3254$\pm$0.08\%  & 2.20$\pm$0.69\%         & 2.30$\pm$0.84\% & 85.43 & 84.99 & 81.59 \\
300	    & 0.3023$\pm$0.09\%  & 3.66$\pm$0.92\%         & 3.96$\pm$0.55\%         & 86.86 & 86.11 & 83.17 \\
400	    & 0.2841$\pm$0.23\%  & 4.84$\pm$0.58\%         & 5.26$\pm$0.83\%         & \textbf{89.71} & 87.01 & 83.45 \\
500	    & 0.2685$\pm$0.26\%  & 5.72$\pm$0.48\%         & 6.21$\pm$0.78\%         & 87.71 & 87.23 & 83.82 \\
600	    & 0.2573$\pm$0.12\%  & 6.20$\pm$0.39\%         & 6.72$\pm$0.18\%         & 88.14 & 85.69 & 83.41 \\
700	    & 0.2499$\pm$0.20\%  & 6.42$\pm$0.49\%         & 6.94$\pm$0.33\%         & 88.00 & \textbf{87.35} & 84.27 \\
800	    & 0.2444$\pm$0.07\%  & 6.54$\pm$0.39\%         & 7.07$\pm$0.15\%         & 88.57 & 86.90 & 84.09 \\
900	    & 0.2407$\pm$0.10\%  & 6.62$\pm$0.31\%         & 7.14$\pm$0.14\%         & 88.57 & 86.56 & \textbf{84.42} \\
1000	& 0.2380$\pm$0.07\%  & 6.65$\pm$0.39\%         & 7.16$\pm$0.10\%         & 89.14 & 87.12 & 83.86 \\
\hline
\end{tabular}
}
\caption{\label{table-emp} Analyses for the code embedding learning process (M=32, K=16). MeanEucDist, NN-cos, and NN-Euc are averaged across 5 runs. ``SNIPS", ``ATIS", and ``TOP" are the test exact match achieved on the three datasets.}
\end{table*}

\subsection{Analysis for Code Convergence}
We study the relationship among a few variables during code learning for the embeddings from pretrained ALBERT-base (Table \ref{table-emp}). During the first 1000 epochs, the mean Euclidean distance between the original and reconstructed embeddings decrease with a decreasing rate. The average number of shared top-20 nearest neighbours according to cosine similarity and Euclidean distances between the two embeddings increase with a decreasing rate. We apply code embeddings trained for different numbers of epochs to ALBERT-base-v2 and finetune on semantic parsing. On SNIPS and ATIS, we find the best validation setting among learning rate \{2,3,4,5,6\}e-5 and weight decay \{0.01, 0.05, 0.01\}. We observe that the test exact match plateaus for code embeddings trained for more than 400 epochs. On Facebook TOP, we use learning rate 2e-5 and weight decay 0.01, and observe the similar trend.

\subsection{Effects of M and K}
We use embeddings from pretrained ALBERT-base-v2 as reference to learn code embeddings with M in \{8, 16, 32, 64\} and K in \{16, 32, 64\}. As shown in Table \ref{table-mk}, after 700 epochs, the MSE loss for embeddings with \textit{larger} M and K converges to \textit{smaller} values in general. With M=64, more epochs are needed for convergence to smaller MSE losses compared to those from smaller M. We apply the embeddings to ALBERT-base-v2 and finetune on SNIPS. In general, larger M yields better performances. Effects of K are less clear when M is large.

\begin{table}[t]
\centering
\tiny
\resizebox{.4\textwidth}{!}{%
\begin{tabular}{cccccc}
\hline \textbf{M} & \textbf{K} & \textbf{epo} & \textbf{MSE} & \textbf{EM}\\ \hline
8 & 16 & 700 & 0.3155$\pm$0.05\% & 85.43 \\
8 & 32 & 700 & 0.3032$\pm$0.04\% & 87.43 \\
8 & 64 & 700 & 0.2944$\pm$0.04\% & 87.43\\
16 & 16 & 700 & 0.2855$\pm$0.05\% & 88.57\\
16 & 32 & 700 & 0.2727$\pm$0.09\% & 88.00\\
16 & 32 & 700 & 0.2669$\pm$0.08\% & 88.14\\
32 & 16 & 700 & 0.2499$\pm$0.20\% & 89.00\\
32 & 32 & 700 & 0.2421$\pm$0.20\% & 89.14\\
32 & 64 & 700 & 0.2396$\pm$0.27\% & 88.29\\
64 & 16 & 700 & 0.2543$\pm$0.47\% & 88.29\\
64 & 16 & 1000 & 0.2256$\pm$1.06\% & 89.71\\
64 & 32 & 700 & 0.2557$\pm$0.37\% & 89.86\\
64 & 32 & 1000 & 0.2159$\pm$0.43\% & 89.71\\
\hline
\end{tabular}
}
\caption{\label{table-mk} Effects of M and K. Mean squared errors (MSE) are averaged over 5 runs. Best validation exact match (EM) is presented for compressed transformer models trained with 0.05 weight decay and \{3,4,5,6,7\}e-5 peak learning rates on SNIPS.}
\end{table}

\section{Conclusion}
Current state-of-the-art task-oriented semantic parsing models are based on pretrained RoBERTa-base (478MB) or BERT-base (420MB). We apply DistilBERT (256MB), ALBERT-large (68MB), and ALBERT-base (45MB), and observe near state-of-the-art performances. We learn compositional code embeddings to compress the model embeddings by 95.15\% $\sim$ 98.46\%, the pretrained encoders by 20.47\% $\sim$ 34.22\%, and observe 97.5\% performance preservation on SNIPS, ATIS, and Facebook TOP. Our compressed ALBERT-large is 54MB and can achieve 99.6\% performances of the previous state-of-the-art models on SNIPS and ATIS. Our technique has potential to be applied to more tasks including machine translation in the future.

\section*{Acknowledgement}
This project is part of the data science industry mentorship program initiated by Andrew McCallum at University of Massachusetts Amherst. We thank the teaching assistants Rajarshi Das and Xiang Lorraine Li for helpful discussion and the instructor Andrew McCallum for valuable feedback. Experiments in this project are conducted on the Gypsum cluster at UMassAmherst. The cluster is purchased with funds from the Massachusetts Technology Collaborative.

\bibliographystyle{acl_natbib}
\bibliography{anthology,emnlp2020}

\end{document}